Solve dx/x(x+y) = dy/-y(x+y) = dz/-(x-y)(2x+2y+z)


1
Expert's answer
2021-05-20T17:28:12-0400

We have 

"\frac{dx}{-x(x+y)} = \frac{dy}{y(x+y)} = \frac{dz}{(x-y)(2x+2y+z)}"

 

"\frac{dx}{-x(x+y)} = \frac{dy}{y(x+y)}\implies \frac{dx}{-x} = \frac{dy}{y}"


by integration, we get "-\ln(x) = \ln(y) - \ln(c_1) \implies c_1 = xy"


Now also 

"\frac{dx}{-x(x+y)} = \frac{dy}{y(x+y)} = \frac{dz}{(x-y)(2x+2y+z)}"



"= \frac{2dx+2dy+dz}{(x-y)z}"

"\implies \frac{dz}{(x-y)(2x+2y+z)} = \frac{2dx+2dy+dz}{(x-y)z}"




"\implies \frac{dz}{2x+2y+z} = \frac{2dx+2dy+dz}{z}"


"\implies zdz = (2x+2y+z) (2dx+2dy+dz)"

By integration, we get


"\frac{z^2}{2} = \frac{(2x+y+z)^2}{2} + \frac{c_2}{2}"




"\implies c_2 = z^2- (2x+y+z)^2"


Hence, solution is :


"c_2 = f(c_1)"




"\implies z^2- (2x+y+z)^2 = f(xy)"

 .